Determine whether the following series converge, converge absolutely, converge conditionally, or diverge.

  1. "\\displaystyle\\sum_{k=1}^\t\u221e" "(\u221ak+1\/\u221ak)"-3
  2. "\\displaystyle\\sum_{k=1}^\t\u221e" "(cos(k\pi/2)/(2k)\pi"
  3. "\\displaystyle\\sum_{k=1}^\t\u221e"(-1)k+1 "(k-1)!(k+1)!\/(2k)!"
  4. "\\displaystyle\\sum_{k=1}^\t\u221e" "(1.3.5...(2k-1))\/(1.4.4.7...(3k-2)))"
1
Expert's answer
2021-10-25T16:23:46-0400

1.

"\int^{\infin}_1\frac{1}{x^p}dx" is convergent if and only if "p>1"


"\displaystyle{\lim_{k\to \infin}}(\sqrt{k}+1/\sqrt k)^{-3}(\sqrt k)^3=1\neq 0"

Since series "\sum(1/k^{3/2})" converge, series "\\displaystyle\\sum_{k=1}^\t\u221e" "(\u221ak+1\/\u221ak)"-3 converge also.


2.

"\\displaystyle\\sum_{k=1}^\t\u221e" "(cos(k\pi/2)/(2k)\pi"="\\displaystyle\\sum_{k=1}^\t\u221e" "\frac{(-1)^{k}\pi}{4k}"


Since "\\displaystyle\\sum_{k=1}^\t\u221e" "\frac{1}{k}" diverge and "\displaystyle{\lim_{k\to \infin}}|(cos(k\pi/2)/(2k)\pi|=0" , the series conditionally convergent.


3.

"\displaystyle{\lim_{k\to \infin}}\frac{(k+2)!k!}{(2k+2)!}\cdot\frac{(2k)!}{(k+1)!(k-1)!}=\displaystyle{\lim_{k\to \infin}}\frac{(k+2)k}{(2k+1)(2k+2)}=1/4<1"


Since "\\displaystyle\\sum_{k=1}^\t\u221e|a_k|" converge, the series absolutely converge.


4.

"\displaystyle{\lim_{k\to \infin}}\frac{a_{k+1}}{a_k}=\displaystyle{\lim_{k\to \infin}}\frac{2(k+1)-1}{3(k+1)-2}=2/3<1"


The series converge.